OpenStudy (zubhanwc3):

given d/dx [ f(2x) ] = f ' (x) and f ' (1) = 1, find f '(2) a) 1/4 b) 1/2 c) 3/4 d) 3/2 e) 5/2

OpenStudy (zubhanwc3):

\[\frac{ d }{ dx }[f(2x)]=f'(x) \] and \[f'(1)=1, \] find \[f'(2)\]

zepdrix (zepdrix):

Take the derivative of the left side, applying the chain rule:\[\Large\rm 2f'(2\color{orangered}{x})=f'(\color{orangered}{x})\]Understand that step?

zepdrix (zepdrix):

Then if you evaluate this at x=1, you would be able to solve for f'(2) without too much trouble! :)\[\Large\rm 2f'(2\cdot\color{orangered}{1})=f'(\color{orangered}{1})\]

OpenStudy (zubhanwc3):

so f'(2) is 2 i presume

zepdrix (zepdrix):

So we have:\[\Large\rm 2f'(2)=f'(1)\]Yes? Can we plug in any information from here? And then solve for f'(2) somehow?

OpenStudy (zubhanwc3):

by doing 2f'(2*2) = f'(2) so it would be 2, im just confirming.

OpenStudy (zubhanwc3):

o, mb, i feel stupid now, its 1/2 because if we divide both sides by 2, f'(2) = f'(1)/2 which is 1/2

zepdrix (zepdrix):

ya there we go! :)